# CoVoMix: Advancing Zero-Shot Speech Generation for Human-like Multi-talker Conversations

### v1

@@ 4.2 Model Configarations
To demonstrate the performance of our methods, the baseline that we compare with is a flow-matching speech synthesis model with phoneme representation, similar to VoiceBox [26] [^3] https://github.com/lucidrains/voicebox-pytorch.git. The baseline contains two models: the acoustic model and the duration model. The acoustic model of baseline is the same as our VoSingle acoustic model, but generates mel-spectrogram from the phoneme sequence. The duration model of baseline is to predict the duration of each phoneme, which is also trained with flow matching objective and has the same architecture with 2 layers and hidden size of 1024.
We train all models from scratch and perform inference on the best performed model on validation set. The text-to-semantic model is trained for 10 epochs, the acoustic model and duration model are trained for 100 epochs. The probability of dropping condition during training is $p_{uncond}=0.3$, and the strength of guidance is $\alpha=0.7$ during inference.
@@ 4.4 Evaluation Metrics
Objective Metrics: We use cosine speaker similarity (SIM), word error rate (WER), Mel cepstral distortion (MCD)  [^4] https://github.com/chenqi008/pymcd, and NISQA  [^5] https://github.com/gabrielmittag/NISQA to evaluate monologue generation results [38]. SIM measures the cosine similarity between speaker embeddings of generated utterance and the acoustic prompt, extracted from WavLM-TDNN [20]. We use Microsoft Speech Recognition API for WER calculation, which measures the correctness and intelligibility. We use an improved MCD metric that adopts the Dynamic Time Warping (DTW) algorithm to find the minimum MCD between two speeches [39]. NISQA [38] measures the speech quality and naturalness of the synthesized speech.
Subjective Metrics: We perform a human evaluation on the generated monologue and dialogue examples. For monologue, we measure naturalness using comparative mean option score (CMOS). We also use the similarity mean option score (SMOS) between the synthesized and prompt speech to measure the speaker similarity. For dialogue, we measure naturalness using comparative mean option score (CMOS). 15 professional linguistic experts provide judges for all subjective evaluations. The instructions of subjective evaluations are provided in Appendix D.
Dialogue Metrics: We assess the naturalness of the generated dialogue speech by 1) Turn-taking Statistics: Utilizing a pre-trained speaker diarization model  [40; 41] [^6] https://github.com/pyannote/pyannote-audio, we measure the duration of inter- and intra-speaker silences, overlapped speech, and active speech. 2) Para-linguistic behaviors: In this study, we only evaluate languter. Employing a laughter detection tool [42] [^7] https://github.com/jrgillick/laughter-detection, we segment instances of laughter and calculate the total count and average duration of laughter events. and 3) Speech consistency: We generate ten dialogues, each with more than five utterances from the target speaker. We randomly select five three-second segments from the target speaker and then compare the cosine similarity of speaker embeddings among these segments.
@@ 5.1 Objective and Subjective metrics
Table 1 shows objective evaluation results for monologue and dialogue generation across various systems.
We observe that the systems leveraging our proposed methods, i.e., CoVoSingle and CoVoMix, achieve higher speaker similarity, lower WER and MCD than baseline on monologue evaluation set. The phoneme-based baseline model requires accurate phoneme-level alignment, however, it is difficult to perform forced alignment using conventional alignment tool [43] [^8] https://github.com/MontrealCorpusTools/Montreal-Forced-Aligner, especially for speech with spontaneous behavior and noisy background. These inaccurate annotations interfere with the model training, and make the model performance severely degraded. By replacing phoneme representation with semantic token sequence, we no longer rely on phoneme-level alignment, thus improving model performance.
Unlike monologue, on the dialogue test set, the ground truth dialogue and CoVoMix have high WER. This is because there is a certain amount of overlapped utterances in the dialogue, but the given transcription is sorted in chronological order of the beginning of each sentence, thus resulting in inconsistencies between transcription and speech in the overlapping parts. Generating each utterance individually and splicing it into a long conversation, such as CoVoSingle, does not have such mismatch, and therefore produces a lower WER. Moverover, constrained by the performance of automatic speaker diarization, we don’t measure speaker similairity for dialogue evaluation set.
In terms of speech quality, we observe that the proposed systems can surpass the ground truth on both monologue and dialogue sets. The flow-matching based acoustic model is able to eliminate background noise, and therefore produces cleaner audio than real data. Since CoVoMix can generate overlapped speech, which may result in a slightly lower NISQA, comparing with CoVoSingle.
Table 1: Objective evaluation results for monologue and dialogue generation across various systems.
Type System SIM $\uparrow$ WER $\downarrow$ MCD $\downarrow$ NISQA $\uparrow$
Monologue GroundTruth 0.59 6.10 / 3.03
Baseline 0.42 15.85 9.45 2.93
CoVoSingle 0.46 9.99 5.03 3.04
CoVoMix 0.46 8.95 5.12 3.01
Dialogue GroundTruth / 14.91 / 2.73
CoVoSingle / 11.77 4.77 2.90
CoVoMix / 19.84 4.25 2.87
The subjective evaluations consistently corroborate the findings of the objective metrics.
@@ 5.2 Dialogue Metrics
We define four turn-taking activities in a dialogue: 1) intra speaker pause (silence between active speech of the same speaker), 2) inter speaker silence (silence between active speech of different speaker), 3) overlapped segments, and 4) active speech of each speaker [14; 8]. Figure 2 shows the distribution of each these turn-taking events, and the similarity to the ground truth reflects the model’s ability to model turn-taking in a dialogue. Despite CoVoSingle can synthesize high-quality monologue, they show poor performance on dialogue turn taking events, especially for intra-speaker pause, inter-speaker silence, and overlap control. Simply concatenating monologue in utterance level and in semantic sequence level results in low variance in inter- and intra- speaker silence distribution, which makes the generation sounds like machine conversation [8]. Compared with ground truth, CoVoMix shows high simlarity in these turn-taking events, making the generated dialogue more human-realistic.
@@
We compute the frequency and duration of spontaneous laughter behaviors across the conversation test set and compared these data across models to observe how similar these statistics are to ground truth. As shown in Figure 3, we notice that all these proposed models are able to generate laughter with frequency similar to ground truth, which shows precise control in these human-like behaviors. Furthermore, CoVoMix generates dialogue with more similar average duration of each laughter compared with ground truth, while CoVoSingle tend to synthesize shorter laughter.
@@
We calculate the speaker similarity between any two pairs of different utterances in a long conversation. Figure 4 shows a heatmap of the cosine similarity between different segments using utterance level concatenation such as CoVoSingle, using semantic sequence level concatenation, and no concatenation such as CoVoMix. A lighter color means that the speakers of the two segments are less similar. As can be seen from the color inconsistencies in the figure, the utterance level concatenation indeed causes dissimilarity in speaker characteristics, especially for non-adjacent utterances. Semantic level concatenation diminishes this speaker drift problem by generating the whole long conversation all at once. Generating the entire semantic sequence and dialogue without concatenation is the best, with the highest speech continuity and consistency.
Figure 4: Speech consistency of different concatenation methods for dialogue generation
6 Conclusion, Limitation and Future Work
@@
We introduced the CoVoMix system for human-like monologue and dialogue generation. The system was composed of an auto-regressive text-to-semantic model and a flow-matching based acoustic model, with semantic token sequence as an intermediate representation. A 2k-hour public phone conversational dataset was leveraged in training the two models of CoVoMix. Through both objective and subjective evaluations, CoVoMix not only achieved high naturalness and zero-shot speaker similarity in monologue and dialogue generations but also demonstrated controllability in dialogue turn-taking and the generation of spontaneous behavior.
@@
We have observed instances of words being omitted or duplicated occasionally in synthesized speech. This was primarily attributed to the text-to-semantic model being an auto-regressive model without forced duration alignment. Additionally, the dataset utilized for this study was sampled at 8 kHz with background noise, factors that contribute to the degradation of speech quality. In future work, we aim to enhance the text-to-semantic model by scaling it up with more data or initializing it with a pre-trained model, and employing super-resolution technology to improve the training data fidelity.
@@ Appendix B Additional Experiments
We perform detailed ablation experiments for model combination, training data, and model size. We also evaluate the model performance in monologue and dialogue voice conversion.
@@ B.1 Ablation Study on Model Combination
Table 2 illustrates our proposed systems in monologue evaluation set, which are a combination of text-to-semantic model and acoustic model. We abbreviate text-to-semantic as T2S and stereo as S.
@@
We first observe that VoMix mono acoustic model has stronger speaker simulation capability than stereo model, because it is more difficult for the latter to predict two long mel-spectrograms simultaneously with independent speaker characteristics. Moreover, we notice that stereo systems show higher NISQA speech quality, indicating that predicting each channel separately causes less distortion in speech quality. Furthermore, when using the same text-to-semantic model, different acoustic models influence WER, indicating that the pronunciation and rhyme also affect the intelligibility of speech. CoVoSinx achieves better speech intelligibility then CoVoSingle due to the use of VoMix.
Table 2: Objective evaluation on monologue for mono and stereo acoustic model
System T2S Acousic Inference SIM WER NISQA
GroundTruth / / / 0.59 6.10 3.03
CoVoSingle CoSingle VoSingle / 0.46 9.99 3.01
CoVoSinx VoMix Mute 2nd seq 0.46 8.78 3.12
CoVoSinx-S VoMix-S Mute 2nd seq 0.36 10.92 3.26
CoVoMix CoMix VoMix / 0.46 8.95 3.01
CoVoMix-S VoMix-S / 0.36 9.46 3.27
B.2 Ablation Study on Data Augmentation for Text-To-Semantic Model
When the total amount of data is deterministic, the diversity of data, such as data duration and the dialogue content, becomes important. Table 3 shows the performance of text-to-semantic model with different data augmentation method on monologue and dialogue generation in terms of correctness. In order for the text-to-semantic model to correctly generate both short and long dialogues, we need to include both short and long sentences in the training speech. Moreover, we can also simulate dialogue given monologues to increase data diversity. which also increase the prediction accuracy of semantic token sequences. After using monologue data of different lengths and synthetic dialog data, the text-to-semantic model achieves the best performance.
@@ B.2 Ablation Study on Data Augmentation for Text-To-Semantic Model
B.3 Ablation Study on Discrete Semantic Representation and Model Size
Figure 8: Speaker similarity across acoustic model of different size
B.4 Monologue and Dialogue Voice Conversion
@@
Table 4 demonstrates the objective results in monologue and dialogue scenario. We notice that in addition to achieving high speaker similarity, these systems can also achieve high spectral similarity, indicating the strong zero-shot voice conversion capability of our proposed system. Moreover, VoMix achieves comparable MCD results in monologue set, and performs better than VoSingle in dialogue set, demonstrating the superiority of the multi-talker acoustic model VoMix for dialogue modeling.
Table 4: Objective evaluation on voice conversion for monologue and dialogue generation
Type System SIM MCD
Monologue VoSingle 0.44 5.64
VoMix 0.46 5.68
Dialogue VoSingle / 8.28
VoMix / 7.68
Appendix C Data Preparation
Algorithm 1 illustrates the dialogue data preparation pipeline, introduced in Section4.1. We slice long dialogues into smaller stereo-channel dialogue with at least two utterances of different speakers, and set the first and last sentences of the processed dialogue to not contain overlaps with other dialogues so as to ensure no extraneous content in the transcription.

### v2

@@ 4.2 Model Configurations
To demonstrate the performance of our methods, the baseline that we compare with is a flow-matching speech synthesis model with phoneme representation, similar to VoiceBox [46]. [^3] https://github.com/lucidrains/voicebox-pytorch (MIT License) The baseline contains two models: the acoustic model and the duration model. The acoustic model of baseline is the same as VoSingle model, but generates mel-spectrogram from the phoneme sequence. The duration model of baseline is to predict the duration of each phoneme, which is also trained with flow matching objective and has the same architecture with 2 layers and hidden size of 1024.
We train all models from scratch and perform inference on the best performed model on validation set. We use 8 NVIDIA TESLA V100 32GB GPUs for training. The text-to-semantic model is trained for 10 epochs with batch size 48. The acoustic model and duration model are trained for 100 epochs with batch size 64. We adopt Adam optimizer with 1e-4 learning rate. The probability of dropping condition during training is $p_{uncond}=0.3$, and the strength of guidance is $\alpha=0.7$ during inference.
@@ 4.4 Evaluation Metrics
Objective Metrics: We use cosine speaker similarity (SIM), word error rate (WER), Mel cepstral distortion (MCD), [^4] https://github.com/chenqi008/pymcd (MIT License) and NISQA [^5] https://github.com/gabrielmittag/NISQA (MIT License) to evaluate generation results [61]. SIM measures the cosine similarity between speaker embeddings of generated utterance and the acoustic prompt, extracted from WavLM-TDNN [62]. We use a market-leading Speech Recognition API for WER calculation, which measures the correctness and intelligibility. We use an improved MCD metric that adopts the Dynamic Time Warping (DTW) algorithm to find the minimum MCD between two speeches [63]. NISQA [61] measures the speech quality and naturalness of the synthesized speech.
Subjective Metrics: We perform a human evaluation on the generated monologue and dialogue examples. For monologue, we measure naturalness using comparative mean option score (CMOS). We also use the similarity mean option score (SMOS) between the synthesized and prompt speech to measure the speaker similarity. For dialogue, we measure naturalness using comparative mean option score (CMOS). 14 professional linguistic experts provide judges for all subjective evaluations. They provide a rating to the second audio, which is randomly selected from a pair of audios, in the (-3 to +3) range. The instructions of subjective evaluations are provided in Appendix F.
Dialogue Metrics: We assess the naturalness of the generated dialogue speech through three metrics: 1) Turn-taking Statistics: By employing a pre-trained speaker diarization model  [64, 65], [^6] https://github.com/pyannote/pyannote-audio (MIT License) we measure the duration of inter- and intra-speaker silences, overlapped speech, and active speech. 2) Para-linguistic Behaviors: Our evaluation focuses on laughter in this study. Employing a laughter detection tool [66], [^7] https://github.com/jrgillick/laughter-detection (MIT License) we identify instances of laughter and calculate both the total count and average duration of these events. and 3) Speech Consistency: To evaluate consistency, we generate ten dialogues, each containing more than five utterances from the target speaker. We then select five three-second segments at random from the target speaker and compare the cosine similarity of speaker embeddings among these segments.
@@ 5.1 Objective and Subjective Metrics
Table 1 shows objective and subjective evaluation results for monologue and dialogue generation across various systems.
We observe that the systems leveraging our proposed methods, i.e., CoVoSingle and CoVoMix, achieve higher speaker similarity, lower WER and MCD than baseline on monologue evaluation set. The phoneme-based baseline model requires accurate phoneme-level alignment, however, it is challenging to perform accurate forced-alignment using conventional alignment tool [67], [^8] https://github.com/MontrealCorpusTools/Montreal-Forced-Aligner (MIT License) especially for speech with spontaneous behavior and noisy background. These inaccuracies in alignment can lead to significant performance degradation. By substituting phoneme representation with semantic token sequences, our approach eliminates the dependency on phoneme-level alignment, thereby enhancing model performance.
The dialogue results show that, unlike monologue, the ground truth and CoVoMix exhibit high WER due to overlapping speech segments. The transcriptions are chronologically sorted, leading to mismatches between transcription and speech in overlapped parts. Furthermore, automatic recognizing overlapped speech while maintaining a low WER remains a challenging task to date. CoVoSingle, which generates utterances separately and combines them, avoids this issue, resulting in lower WER.
In terms of speech quality, we observe that the proposed systems can surpass the ground truth on both monologue and dialogue sets. The flow-matching based acoustic model is able to eliminate background noise, and therefore produces cleaner audio than real data. CoVoMix can generate overlapped speech, which may result in a slightly lower NISQA, comparing with CoVoSingle.
Table 1: Objective and subjective evaluation results for monologue and dialogue generation across various systems.The symbol $\dagger$" is used to indicate that the system performance is significantly different (p<0.01) from CoVoSingle system in terms of CMOS and SMOS scores.
Eval Set System SIM $\uparrow$ WER $\downarrow$ MCD $\downarrow$ NISQA $\uparrow$ CMOS $\uparrow$ SMOS $\uparrow$
Monologue GroundTruth 0.59 6.10 / 3.03 / /
Baseline 0.42 15.85 9.45 2.93 -1.60$\dagger$ -1.18$\dagger$
CoVoSingle 0.49 9.99 6.15 3.04 0.00 0.00
CoVoMix 0.49 8.95 6.04 3.01 0.83$\dagger$ 0.11
Dialogue GroundTruth / 14.91 / 2.73 / /
CoVoSingle / 11.77 6.91 2.90 0.00 /
CoVoMix / 19.84 6.82 2.87 0.81$\dagger$ /
The subjective evaluations consistently support the findings of the objective metrics. As shown in Table 1, CoVoSingle significantly outperforms baseline in terms of both CMOS and SMOS scores for monologue testing set. Furthermore, across both monologue and dialogue testing sets, CoVoMix demonstrates signifcantly better performance over CoVoSingle on CMOS scores.
We have not found a good way to measure the similarity for the dialogue testing set due to the necessity of speaker diarization. The potential errors in speaker diarization could impact the fairness of the comparison. Additionally, it poses a challenge for a subject to assign a SMOS score in scenarios involving multiple talkers.
@@ 5.2 Dialogue Metrics
We define four turn-taking activities in a dialogue: 1) intra speaker pause (silence between active speech of the same speaker), 2) inter speaker silence (silence between active speech of different speaker), 3) overlapped segments, and 4) active speech of each speaker [15, 8]. Figure 3 shows the distribution of various turn-taking activities. The degree of similarity to the ground truth reflects the model’s ability to simulate turn-taking in a dialogue. While CoVoSingle can synthesize high-quality monologue, it exhibits subpar performance in dialogue turn taking events, particularly in managing intra-speaker pause, inter-speaker silence, and overlap control. Simply concatenating monologue at utterance level results in low variance in inter- and intra- speaker silence distribution, leading in a dialogue which sounds robotic and lacks the natural flow of conversation [8]. In contrast, CoVoMix demonstrates a high similarity to the ground truth in these turn-taking events, yielding more human-realistic dialogues.
@@
We computed the frequency and duration of spontaneous laughter behaviors across the conversation test set and compared these metrics across models to check their closeness to the ground truth. As illustrated in Figure 4, it shows that all proposed models can generate laughter with a frequency close to the ground truth, demonstrating precise control over these human-like behaviors. Moreover, CoVoMix can produce dialogues with an average laughter duration that is closer to the ground truth, whereas CoVoSingle tends to synthesize shorter instances of laughter.
@@
We calculate the speaker similarity between any two pairs of different utterances in a long conversation. Figure 5 presents a heatmap of the cosine similarity between different segments, contrasting utterance-level concatenation methods like CoVoSingle with non-concatenation approaches like CoVoMix. A lighter shade indicates lower speaker similarity. The figure’s color inconsistencies reveal that utterance-level concatenation can indeed lead to dissimilar speaker characteristics, particularly for non-adjacent utterances. Generating the entire dialogue without concatenation results in significantly improved consistency of speaker similarity across various utterances.
Figure 5: Speech consistency of CoVoSingle and CoVoMix for dialogue generation
6 Ablation Studies and Extension
To enhance the effectiveness of text-to-semantic modeling, we conducted ablation studies focusing on data augmentation and model size. In addition to real dialogue data, we incorporated simulated dialogues and monologue sentences into training data. Results show the benefits of such augmentation, as evidenced by improved model prediction accuracy, i.e., reduced WER, in both monologue and dialogue generation tasks. Furthermore, we explored the impact of output channel configurations for the acoustic model by comparing single-channel mixed speech output with dual-channel outputs, where each channel contained speech from an individual speaker. Experimental results show that dual-channel outputs underperformed in WER, and outperformed in NISQA compared with single-channel outputs. Please refer to Appendix B for the detailed results of all ablation studies.
Our acoustic model can generate specific speakers’ voices, given semantic token sequences and target speakers’ prompts. So it is straightforward to be extended to a voice conversion task, which modifies the speech of a source speaker and makes their speech sound like that of another target speaker without changing the content information. Instead of predicting semantic tokens from given text, we extract the semantic tokens from the speech of the source speaker. VoSingle performs voice conversion of dialogue by processing each channel individually and then mix them up, while VoMix model achieves voice conversion simultaneously. We notice that in addition to achieving high speaker similarity, these systems can also achieve high spectral similarity, indicating the strong zero-shot voice conversion capability. Moreover, VoMix performs better than VoSingle in both monologue and dialogue sets. The detailed results are shown in Appendix C and the corresponding demo is provided in https://aka.ms/covomix.
7 Conclusion, Limitation, Future Work and Broader Impacts
@@
We introduce the CoVoMix system for human-like monologue and dialogue generation. The system is composed of an auto-regressive text-to-semantic model and a flow-matching based acoustic model, with semantic token sequence as an intermediate representation. A 2k-hour conversational telephone speech dataset is leveraged in training these two models of CoVoMix. Through both objective and subjective evaluations, CoVoMix not only achieves high naturalness and zero-shot speaker similarity in both monologue and dialogue generations but also demonstrates its proficiency in the fluency of dialogue turn-taking and spontaneous behavior generation.
@@
Limitation and Future work We have observed instances of words being omitted or duplicated occasionally in synthesized speech. This is primarily attributed to the text-to-semantic model being an auto-regressive model without forced duration. Additionally, the dataset utilized for this study is sampled at 8 kHz with background noise, factors that contribute to the degradation of speech quality. In future work, we aim to enhance the text-to-semantic model by scaling it up or initializing it with a pre-trained model, and employing super-resolution methods to improve the training data fidelity.
Broader Impacts A high-quality and human-like speech generation model like CoVoMix can enable many applications that improve the quality of our life. However, since CoVoMix could synthesize speech that maintains speaker identity, it may carry potential risks in misuse of the model, such as spoofing voice identification or impersonating a specific speaker. To mitigate such risks, it is possible to build a detection model to discriminate whether an audio clip was synthesized by CoVoMix.
@@ Appendix B Additional Experiments
We perform detailed ablation experiments for model combination, model size and training data.
@@ B.1 Ablation Study on Model Combination
Table 2 illustrates our proposed systems in monologue and dialogue evaluation set, which are a combination of text-to-semantic model and acoustic model. We abbreviate text-to-semantic as T2S and stereo as S.
@@
We first observe that when using the same text-to-semantic model, different acoustic models influence WER, indicating that the pronunciation and rhyme also affect the intelligibility of speech. For example, CoVoSinx achieves better speech intelligibility than CoVoSingle due to the use of VoMix. The speaker similarity performance for monologue generation is the similar across models.
For dialogue generation, we notice that stereo systems show comparable WER than mono systems. Moreover, we observe that stereo systems show higher NISQA speech quality, indicating that predicting each channel separately causes less distortion in speech quality than predicting mixed mel-spectrogram.
Table 2: Objective evaluation on monologue and dialogue for mono and stereo acoustic model
System T2S Acousic Monologue Dialogue
SIM $\uparrow$ WER $\downarrow$ NISQA $\uparrow$ WER $\downarrow$ NISQA $\uparrow$
GroundTruth / / 0.59 6.10 3.03 14.91 2.73
CoVoSingle CoSingle VoSingle 0.49 9.99 3.01 11.76 2.90
CoVoSinx VoMix 0.49 8.78 3.12 12.27 2.97
CoVoSinx-S VoMix-S / / / 12.95 3.19
CoVoMix CoMix VoMix 0.49 8.95 3.01 19.84 2.87
CoVoMix-S VoMix-S / / / 20.35 3.00
B.2 Ablation Study on Discrete Semantic Representation and Model Size
Figure 9: Speaker similarity across acoustic model of different size
B.3 Ablation Study on Data Augmentation for Text-To-Semantic Model
When the total amount of data is deterministic, the diversity of data, such as data duration and the dialogue content, becomes important. Table 3 shows the performance of text-to-semantic model with different data augmentation method on monologue and dialogue generation in terms of correctness. In this study, the acoustic model used only contains 2 transformer layers instead of 8 layers mentioned in 4.2 for faster inference. We incorporate a diverse training dataset including both short and long sentences for the text-to-semantic model to accurately generate dialogues of varying lengths. The long monologue has a minimum duration of 10 seconds, whereas the short monologue has minimum duration of 1 second. Additionally, we enhance data variety by simulating dialogues from monologues through concatenation, which also improved the prediction accuracy of semantic token sequences. As a result of utilizing monologue data of different lengths and synthetic dialogue data, the text-to-semantic model demonstrated the best performance.
@@ B.3 Ablation Study on Data Augmentation for Text-To-Semantic Model
Appendix C Extension to Voice Conversion
@@
Table 4 demonstrates the objective results in monologue and dialogue scenario. We notice that in addition to achieving high speaker similarity, these systems can also achieve high spectral similarity, indicating the strong zero-shot voice conversion capability of our proposed system. Moreover, VoMix performs better than VoSingle in both monologue and dialogue sets.
Table 4: Objective evaluation on voice conversion for monologue and dialogue generation. The symbol $\dagger$" is used to indicate that the system performance is significantly different (p<0.01) from VoSingle system
Type System SIM $\uparrow$ MCD $\downarrow$
Monologue VoSingle 0.47 6.47
VoMix 0.49$\dagger$ 6.46
Dialogue VoSingle / 6.70
VoMix / 6.59$\dagger$
Appendix D Experiment Statistical Significance
In order to determine statistical significance of the main experiment in Table 1, we first use numpy.mean and numpy.std in python to calculate the mean and the standard deviation of objective metrics across the test set in Table 5. Moreover, we use z-test to determine if the differences are statistically significant. We notice that results are statistically significant for WER, MCD and NISQA in both monologue and dialogue evaluation sets. Similar to subjective evaluation result, the speaker similarity performance of the CoVoSingle and CoVoMix systems are relatively close and do not show significant differences. Besides, the WER has a large deviation because the text-to-semantic model might synthesize speech with omitted or duplicated words, as mentioned in the limitation part in Section 7.
Table 5: Objective evaluation results for monologue and dialogue generation across various systems. The symbol $\dagger$" is used to indicate that the system performance is significantly different (p<0.01) from CoVoSingle system
Eval Set System SIM $\uparrow$ WER $\downarrow$ MCD $\downarrow$ NISQA $\uparrow$
Monologue CoVoSingle 0.49±0.17 9.99±9.02 6.15±1.85 3.04±0.39
CoVoMix 0.49±0.18 8.95±8.68$\dagger$ 6.04±2.03$\dagger$ 3.01±0.44$\dagger$
Dialogue CoVoSingle / 11.77±9.00 6.91±1.87 2.90±0.28
CoVoMix / 19.84±19.83$\dagger$ 6.82±2.12$\dagger$ 2.87±0.37$\dagger$
To investigate the effect of randomness, we assess the same model using three different random seeds. For each seed, we evaluate the model performance and compute the mean and standard deviation across different seeds. The results are shown in Table 6. Our findings indicate that the standard deviation among the different random seeds is relatively small, suggesting that the system exhibits stability in the presence of randomness.
Table 6: Objective evaluation results for monologue and dialogue generation across various systems with different random seeds
Eval Set System SIM $\uparrow$ WER $\downarrow$ MCD $\downarrow$ NISQA $\uparrow$
Monologue CoVoSingle 0.484±0.005 10.206±0.166 6.147±0.017 3.035±0.014
CoVoMix 0.488±0.005 9.378±0.349 6.058±0.015 3.008±0.002
Dialogue CoVoSingle / 11.903±0.262 6.916±0.031 2.902±0.002
CoVoMix / 19.542±0.587 6.829±0.006 2.870±0.008
Appendix E Data Preparation
Algorithm 1 illustrates the dialogue data preparation pipeline, introduced in Section4.1. The hyper-parameter maxDuration is set to 40 seconds by default.